Find the present equivalent value of a series of cash flows. We expect an $5,103 series of positive cash flows annually, and in addition negative cash flows of 5,524 in 3rd year and 10,908 in 5th year of operation. Assume an interest rate of 0.09, and 8 years of operation.
Answer #1

We fnd the PV by multiplying each year cash flow with Discount Factor . Discount factor for each year can be calculated as 1/ (1+r)^N

For year 1 : 1/(1+9%)^1 = 0.92

Cashflows for year 3 = 5103 - 5524 = -421

Cashflows for year 5 = 5103 - 10908 = -5805

9%
Year Cashflows DF PV
1 5103                         0.92 4681.7
2 5103                         0.84 4295.1
3 -421                         0.77 -325.1
4 5103                         0.71 3615.1
5 -5805                         0.65 -3773
6 5103                         0.60 3042.8
7 5103                         0.55 2791.5
8 5103                         0.50 2561
Total 16889

PV of cash flows is 16,889

> how to find the value of PV?

jagadeesh varma Wed, Apr 14, 2021 10:27 AM